Compute the corresponding (OH-) or (H3O+) in each of the following conditions. Describe the condition as acidic or base.


1. (H3O+) = 2.3 x 10-9 M

2. (OH-) = 7.4 x 10-3 M

3. (H3O+) = 8.7 x 10-4 M



1
Expert's answer
2021-04-13T08:04:26-0400

1. PH = -log (H3O+) = -log 2.3 x 10-9 M = 8.64 which is basic

2. POH = -log (OH-) = -log 7.4 x 10-3 M = 2.13 which is basic

3. PH = -log (H3O+) = -log 8.7 x 10-4 M = 3.06 which is acidic